OpenStudy (anonymous):

y' - 2y = e^3t, y(0) = 3, Find the general solution and solve for the IVP. I have no idea how to solve non homogeneous differential equations

OpenStudy (jackellyn):

To solve non homogeneous you need to find the general solution of the homogeneous part y'-2y=0 which would be Yc and then the particular solution (Yp) for the right hand side =e^3t the general solution would be y=Yc+Yp

OpenStudy (anonymous):

The correct answer is e^3t + 2e^2t. I understand how to come up with this but when solving for the homogeneous part I get y=(e^(t^2)/2)/2. What am I doing wrong here?

OpenStudy (jackellyn):

If you use the characteristic equation, how would your answer be?

OpenStudy (jackellyn):

r-2=0 right?

OpenStudy (anonymous):

I assumed that y' = 2y(t) and then got dy/2y =1dt. DOes the 2 go with the t rather then the y

OpenStudy (anonymous):

I have never been told to use r as anything before

OpenStudy (jackellyn):

You can get a step by step soln there

OpenStudy (anonymous):

Okay thank you!

OpenStudy (anonymous):

Also for 2nd order how would you solve y"-y'=0.

OpenStudy (anonymous):

nevermind sorry I'll just use that website

OpenStudy (jackellyn):

I learned using characteristic equations where you substitute y''=r^2, y'=r and then you factor which will give you zeros y''-y'=0 r^2-r=0 r(r-1)=0 r=0 r=1 general soln \[y=C _{1}+C _{2}e ^{x}\]